How is w=(3,5,1)∈R

3

 a linear combination of u=(0,−2,2)

 and v=(1,3,−1)

 ?


1
Expert's answer
2022-04-11T16:48:47-0400

w is a linear combination of u and v if exist such x and y so x*u +y*v = w

That means that "\begin{bmatrix}
 0 & 1 \\
 -2 & 3 \\
 2 & -1
\end{bmatrix} \cdot
\begin{bmatrix}
 x \\
 y \\
\end{bmatrix} =
\begin{bmatrix}
 3 \\
 5 \\
 1
\end{bmatrix}"


Or "\begin{cases}
 y = 3 \\
 -2 x + 3y = 5 \\
 2x - y = 1
\end{cases}"

From the fist equation find that y = 3, and substitute it in the last equation get 2x -3 = 1, or x = 2. After substituting x=2 and y=3 in the second equation confirmed it is satisfied.

So w = 2*u + 3*v